# A General Framework for Inference-time Scaling and Steering of Diffusion Models

### v1

@@ 1 Introduction
We introduce Feynman-Kac steering (fk steering), an inference-time steering framework for sampling tilted distributions with diffusion models. fk steering enables guidance with arbitrary reward functions, differentiable or otherwise, for both discrete and continuous-state space models. fk steering makes use of a rare-event simulation method, fk interacting particle systems (Moral, 2004; Hairer and Weare, 2014; Vestal et al., 2008). This technique enables the generation of samples with high-rewards, which may be rare events under the original model $p_{\theta}(\mathbf{x}_{0})$.
@@ 2 Related Work
fk steering builds on top of recent work such as tds (Wu et al., 2023b) and others (Trippe et al., 2022; Cardoso et al., 2023) that propose inference-time techniques, which make use of particle-based sampling methods for sampling from the unnormalized target distributions.
@@
More recently, Li et al. (2024) propose svdd, a derivative-free approach to guiding diffusion models for reward maximization, by targeting the distribution the limit of $\lim_{\lambda\rightarrow\infty}\frac{1}{\mathbf{Z}}p_{\theta}(\mathbf{x}_{0})\exp(\lambda r(\mathbf{x}_{0}))$. If there is a single sample $\mathbf{x}^{*}_{0}$ in the support of $p_{\theta}$ that maximizes $r$, then this distribution collapses to a point mass on $\mathbf{x}^{*}_{0}$. At each time step, svdd samples $k$ states $\mathbf{x}_{t}^{i}$ from a diffusion model and selects one state $\mathbf{x}_{t}^{i*}$ with the highest reward and makes $k$ copies of that state, therefore reducing diversity of samples $\mathbf{x}_{0}^{i}$.
@@ 3.3 Feynman-Kac diffusion steering
We use fk-ips to produce paths $(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{0})$ with high-reward $\mathbf{x}_{0}$ samples. fk-ips tilts each transition kernel $p_{\theta}(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1})$ with a potential function $G_{t}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t})$:
$p_{\textsc{fk}}(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1},\dots,\mathbf{x}_{T},\mathbf{c})$ $\propto p_{\theta}(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1},\mathbf{c})G_{t}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t},\mathbf{c})$ (3)
which defines the following distribution over paths $p_{\textsc{fk}}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t})$:
$p_{\textsc{fk}}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t}\,|\,\mathbf{c})=\frac{1}{\mathbf{Z}_{t}}p_{\theta}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t}\,|\,\mathbf{c})\left\{\prod_{s}G_{t}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{s},\mathbf{c})\right\}$ (4)
@@
This choice ensures that $p_{\textsc{fk}}(\mathbf{x}_{T},\dots,\mathbf{x}_{0})\propto p_{\theta}(\mathbf{x}_{T},\dots,\mathbf{x}_{0}\,|\,\mathbf{c})\exp(\lambda r(\mathbf{x}_{0},\mathbf{c}))$; that is, sampling from $\mathbf{x}_{0}\sim p_{\textsc{fk}}$ samples from the target that tilts towards higher rewards. Potential functions that satisfy this constraint are not unique.
Sampling from $p_{\textsc{fk}}$.
For most choices of the potential $G_{t}$, sampling from the fk distribution in eq. 3 is intractable. In practice, fk steering makes use of particle-based approximations such as ticketed diffusion Monte Carlo (dmc) (Hairer and Weare, 2014), residual resampling (Liu and Liu, 2001), or smc (Moral, 2004). All these methods have three ingredients in common (1) sampling multiple particles, (2) scoring the particles using the potentials, and (3) duplicating high-scoring particles and terminating low-scoring particles. For algorithms such as dmc, the number of particles during generation is random. Alternatively, in smc the number of particles is fixed.
smc generates $k$ particles using a proposal generator $\tau(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1},\dots,\mathbf{x}_{T},\mathbf{c})$ and at each transition step $\mathbf{x}_{t+1}\rightarrow\mathbf{x}_{t}$ scores the particles using the potential and the transition kernel importance weights:
$G^{i}_{t}=G_{t}(\mathbf{x}^{i}_{T},\dots,\mathbf{x}_{t+1}^{i},\mathbf{x}_{t}^{i},\mathbf{c})\frac{p_{\theta}(\mathbf{x}^{i}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}^{i}_{t+1},\mathbf{c})}{\tau(\mathbf{x}^{i}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}^{i}_{t+1},\dots,\mathbf{x}^{i}_{T},\mathbf{c})}.$ (6)
@@
The ideal rewards for the intermediate state $\mathbf{x}_{t}$ would require knowing the distribution of rewards of $\mathbf{x}_{0}$ generated from the state $\mathbf{x}_{t}$: $p_{\theta}(r(\mathbf{x}_{0})\,|\,\mathbf{x}_{t},\mathbf{c})$. With this distribution, reward functions $r_{\phi}$ could be chosen to yield high expected reward when reward variance is low, or based on the 10th percentile of the reward distribution to ensure good worst-case quality. Producing this distribution of rewards requires training with samples from the model. We will describe a few options that have different tradeoffs in compute versus knowledge provided about the rewards for the sample $\mathbf{x}_{0}$.
@@ C.1 Choice of proposal distribution
    and sample from $p_{\textsc{fk}}(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1},\dots,\mathbf{x}_{T})\propto p_{\theta}(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1})G_{t}(\mathbf{x}_{T},\dots,\mathbf{x}_{t})$.
@@ C.2 How existing work fits into fk steering
With $\lambda=\infty$, svdd is equivalent to doing best-of-$n$ at each step, since the authors recommend sampling from $p_{\text{target}}(\mathbf{x}_{0})\propto\lim_{\lambda\rightarrow\infty}p_{\theta}(\mathbf{x}_{0})\exp(\lambda r(\mathbf{x}_{0}))$. We note that similar to svdd, $p_{\textsc{fk}}$ can be sampled using nested importance sampling.

### v3

@@ 1 Introduction
We introduce Feynman-Kac steering (fk steering), an inference-time steering framework for sampling tilted distributions with diffusion models. fk steering enables guidance with arbitrary reward functions, differentiable or otherwise, for both discrete and continuous-state space models. fk steering makes use of a rare-event simulation method, fk interacting particle systems (Moral, 2004; Hairer and Weare, 2014; Vestal et al., 2008; Chopin et al., 2020). This technique enables the generation of samples with high-rewards, which may be rare events under the original model $p_{\theta}(\mathbf{x}_{0})$.
@@ 2 Related Work
fk steering builds on top of recent work such as tds (Wu et al., 2023b) and others (Trippe et al., 2022; Cardoso et al., 2023; Dou and Song, 2024) that propose inference-time techniques, which make use of particle-based sampling methods for sampling from the unnormalized target distributions.
@@
More recently, Li et al. (2024) propose svdd, a derivative-free approach to guiding diffusion models for reward maximization, by targeting the distribution the limit of $\lim_{\lambda\rightarrow\infty}\frac{1}{\mathbf{Z}}p_{\theta}(\mathbf{x}_{0})\exp(\lambda r(\mathbf{x}_{0}))$. If there is a single sample $\mathbf{x}^{*}_{0}$ in the support of $p_{\theta}$ that maximizes $r$, then this distribution collapses to a point mass on $\mathbf{x}^{*}_{0}$. svdd uses nested importance sampling for sampling, see algorithm 5 in Naesseth et al. (2019). At each time step, svdd samples $k$ states $\mathbf{x}_{t}^{i}$ from a diffusion model and selects one state $\mathbf{x}_{t}^{i*}$ with the highest reward and makes $k$ copies of that state, therefore reducing diversity of samples $\mathbf{x}_{0}^{i}$.
@@ 3.3 Feynman-Kac diffusion steering
We use fk-ips to produce paths $(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{0})$ with high-reward $\mathbf{x}_{0}$ samples. fk-ips defines a sequence of distributions $p_{\textsc{fk},t}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t})$ by tilting the base distribution $p_{\theta}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t})$ using potentials $G_{t}$ (Moral, 2004; Chopin et al., 2020). The sequence of distributions $p_{\textsc{fk},t}$ is built iteratively by tilting the transition kernels $p_{\theta}(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1})$ with a potential $G_{t}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t})$. We start with $p_{\textsc{fk},T}(\mathbf{x}_{T})\propto p_{\theta}(\mathbf{x}_{T}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{c})G_{T}(\mathbf{x}_{T},\mathbf{c})$ and then define the subsequent distributions as:
$p_{\textsc{fk},t}(\mathbf{x}_{T},\dots,\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{c})$ $=\frac{1}{\mathbf{Z}_{t}}p_{\theta}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{c})\left\{\prod_{s=T}^{t}G_{t}(\mathbf{x}_{T},\mathbf{x}_{T-1},\dots,\mathbf{x}_{s},\mathbf{c})\right\}$ (3)
$=\frac{1}{\mathbf{Z}_{t}}\underbrace{p_{\theta}(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1},\mathbf{c})G_{t}(\mathbf{x}_{T},\dots,\mathbf{x}_{t},\mathbf{c})}_{\text{Tilted transition kernel}}\underbrace{p_{\theta}(\mathbf{x}_{T},\dots,\mathbf{x}_{t+1}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{c})\prod_{s=T}^{t+1}G_{s}(\mathbf{x}_{T},\dots,\mathbf{x}_{s},\mathbf{c})}_{\propto\text{Previous }p_{\textsc{fk},t+1}}$ (4)
@@
This choice ensures that $p_{\textsc{fk},0}(\mathbf{x}_{T},\dots,\mathbf{x}_{0}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{c})\propto p_{\theta}(\mathbf{x}_{T},\dots,\mathbf{x}_{0}\,|\,\mathbf{c})\exp(\lambda r(\mathbf{x}_{0},\mathbf{c}))$; that is, sampling from $\mathbf{x}_{0}\sim p_{\textsc{fk},0}$ samples from the target that tilts towards higher rewards. Potential functions that satisfy this constraint are not unique.
Sampling from $p_{\textsc{fk},0}$.
Direct sampling from $p_{\textsc{fk},0}$ is intractable. However, targeting the intermediate distributions $p_{\textsc{fk},t}$ supports sampling of the distribution $p_{\textsc{fk},0}$ through the use of particle-based methods, such as smc (Moral, 2004; Doucet and Lee, 2018), nested is (see alg. 5 in Naesseth et al. (2019)), diffusion Monte Carlo (dmc) (Hairer and Weare, 2014). smc generates $k$ particles using a proposal generator $\tau(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1},\dots,\mathbf{x}_{T},\mathbf{c})$ and at each transition step $\mathbf{x}_{t+1}\rightarrow\mathbf{x}_{t}$ scores the particles using the potential and the transition kernel importance weights:
$G^{i}_{t}$ $=\frac{p_{\textsc{fk},t}(\mathbf{x}_{T},\dots,\mathbf{x}_{t+1},\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{c})}{p_{\textsc{fk},t+1}(\mathbf{x}_{T},\dots,\mathbf{x}_{t+1}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{c})\tau(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1},\dots,\mathbf{x}_{T}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{c})}$ (6)
$=G_{t}(\mathbf{x}^{i}_{T},\dots,\mathbf{x}_{t+1}^{i},\mathbf{x}_{t}^{i},\mathbf{c})\frac{p_{\theta}(\mathbf{x}^{i}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}^{i}_{t+1},\mathbf{c})}{\tau(\mathbf{x}^{i}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}^{i}_{t+1},\dots,\mathbf{x}^{i}_{T},\mathbf{c})}.$ (7)
@@
The ideal rewards for the intermediate state $\mathbf{x}_{t}$ would require knowing the distribution of rewards of $\mathbf{x}_{0}$ generated from the state $\mathbf{x}_{t}$: $p_{\theta}(r(\mathbf{x}_{0})\,|\,\mathbf{x}_{t},\mathbf{c})$. With this distribution, reward functions $r_{\phi}$ could be chosen to yield high expected reward when reward variance is low, or based on the 10th percentile of the reward distribution to ensure good worst-case quality. Producing this distribution of rewards requires training with samples from the model. We will describe a few options that have different trade-offs in compute versus knowledge provided about the rewards for the sample $\mathbf{x}_{0}$.
@@ C.1 Choice of proposal distribution
    and sample from $p_{\textsc{fk},(}\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1},\dots,\mathbf{x}_{T})\propto p_{\theta}(\mathbf{x}_{t}\leavevmode\nobreak\ |\leavevmode\nobreak\ \mathbf{x}_{t+1})G_{t}(\mathbf{x}_{T},\dots,\mathbf{x}_{t})$.
@@ C.2 How existing work fits into fk steering
With $\lambda=\infty$, svdd is equivalent to doing best-of-$n$ at each step, since the authors recommend sampling from $p_{\text{target}}(\mathbf{x}_{0})\propto\lim_{\lambda\rightarrow\infty}p_{\theta}(\mathbf{x}_{0})\exp(\lambda r(\mathbf{x}_{0}))$. We note that similar to svdd, $p_{\textsc{fk},0}$ can be sampled using nested importance sampling.